kekeman:

Find the exact value of tan5pie/3. A.) √(3) B.) -√(3) C.) -1 D.) √(3)/2

Narad:

From the trigonometric circle

1 attachment
Narad:

\[\tan \left( \frac{ 5\pi }{ 3} \right)=-\tan \left( \frac{ \pi }{ 3 } \right)\]

kekeman:

@narad wrote:
\[\tan \left( \frac{ 5\pi }{ 3} \right)=-\tan \left( \frac{ \pi }{ 3 } \right)\]
I think it is -√(3)

kekeman:

Am I correct?

Narad:

Yes, that's the correct answer